Determine whether (-4, -12) is a solution of the system. {x-y=8
{2x+y=20

Answer by stanbon(48558)

Substute -4 in place of x and -12 in place of y and see if you
get a true result on both equations.
-4-(-12)=8
8=8
2(-4)+(-12)= 2o
-8-12 =20
-20=20
The point gives a true result in the first equation but not in the second one.
The point (-4,-12) is not a solution for this system of equations.
Cheers,
Stan H.